OpenStudy (zeesbrat3):

Find the values of m and b that make the following function differentiable. the piecewise function f(x) = x^3 when x is less than or equal to 1 or mx + b when x is greater than one

OpenStudy (zeesbrat3):

\[f(x) = \left\{ \left[\begin{matrix}x^3, & x \le1 \\ mx+b & x >1\end{matrix}\right] \right\}\]